MuleSoft, Inc. conducted its IP on March 17, 2017 for the principle purposes of increasing its capitalization and financial flexibility, creating a public market for its Class A common stock, and enabling access to the public equity markets for it and its stockholders. MuleSoft sold 13 million shares for an IPO offer of $17 per share. The underwriting discoung was $1.19 per share. MuleSoft intends to use the net proceeds from the offering to the firm for general corporate purposes, such as working capital, operating expenses, capital expenditures, and to possibly acquire complementary businesses, products, services or technologies. MuleSoft's closing stock price was $24.75 after the first day of trading on the NYSE and there were 125,991,557 shares of stock outstanding.

Create a spreadsheet to conduct an analysis of MuleSoft's IPO and determine the following:

a) Calculate the total proceeds for MuleSoft's IPO.

b) Calculate the percentage underwriter discount for MuleSoft's IPO.

c) Calculate the dollar amount of the underwriting fee for MuleSoft's IPO.

d) calculate the net proceeds for MuleSoft's IPO.

e) Calculate the percentage IPO underpricing for MuleSoft's IPO.

f) Calculate the market capitalization for MuleSoft's IPO after the first day of trading in the secondary market.
Answer #1

Please note that a spreadsheet can't be attached along with the solution. Hence, I am producing below a snap shot from my spreadsheet. Please look at it step by step. In each step, i have explained how a particular calculation has bee done. Please use this to create your own spreadsheet.

Part of the question

Particular

Unit

Value

Numbers of shares issued in IPO, N

million

13.00

IPO Price per share, P

$

17.00

(a)

Total proceeds for MuleSoft's IPO, S

$ million

221.00

(Calculated as P x N = 13 x 17)

Underwriting discount per share, D

$

1.19

(b)

percentage underwriter discount for MuleSoft's IPO

%

7.00%

(Calculated as D / P = 1.19 / 17)

(c )

Dollar amount of the underwriting fee for MuleSoft's IPO, F

$ million

15.47

(Calculated as Discount per share x Nos. of shares issued = D x N = 1.19 x 13)

(d)

Net proceeds for MuleSoft's IPO

$ million

205.53

(Calculated as Total proceeds from IPO - Dollar amount of underwriting fees = S - F = 221 - 15.47)

Closing price on 1st day of trading, P1

$

24.75

Let's assume market discovered the correct price of the stock on the first day of trading

Underpricing in the IPO, L

$

7.75

(Calculated as 1st trading day close price - IPO price = P1 - P = 24.75 - 17

(e )

Percentage IPO underpricing for MuleSoft's IPO

31.31%

(Calculated as expressing underpricing as %age of tst trading day close price = L / P1 = 7.75 / 24.75

Nos. of shares outstanding at the end of 1st trading day, N1

nos.

125,991,557

(g)

Market capitalization for MuleSoft's IPO after the first day of trading in the secondary market

$

3,118,291,036

(Calculated as Close price x nos. of shares outstanding = P1 x N1 = 24.75 x 125,991,557)